# Deepfake Media Forensics: State of the Art and Challenges Ahead

### v1

@@ VI-C Trusted Remote Media Processing on Cloud and Edge Computing Systems
VII Conclusion
In this paper, we have conducted a comprehensive review of the state-of-the-art techniques and challenges in Deepfake media forensics. Our exploration covered the core areas of Deepfake detection, attribution and recognition, passive authentication, detection in realistic scenarios, and active authentication. Each of these areas addresses specific facets of the Deepfake phenomenon, from the identification of synthetic media and tracing their origins to ensuring the robustness of detection systems in real-world environments and embedding verifiable information within media for instant authentication. Our review underscores the significant progress made in the field and we also acknowledge the ongoing challenges, particularly in addressing the ethical and security risks posed by the misuse of Deepfake technology. Future work will focus on conducting a more in-depth analysis of practical countermeasures and gaining deeper insights into real-world applications (e.g. highly compressed data). Furthermore, a thorough study on the generalization capability of state-of-the-art deepfake detectors will be undertaken. In addition, we will explore issues related to scalability, real-time processing capabilities and ease of use of state-of-the-art deepfake detector approaches.
Acknowledgment
This study has been partially supported by SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU

### v2

@@ 6.3 Trusted Remote Media Processing on Cloud and Edge Computing Systems
7 Conclusion
In this paper, we have conducted a comprehensive review of the state-of-the-art techniques and challenges in Deepfake media forensics. Our exploration covered the core areas of Deepfake detection, attribution and recognition, passive authentication, detection in realistic scenarios, and active authentication. Each of these areas addresses specific facets of the Deepfake phenomenon, from the identification of synthetic media and tracing their origins to ensuring the robustness of detection systems in real-world environments and embedding verifiable information within media for instant authentication. Future work will focus on conducting a more in-depth analysis of practical countermeasures and gaining deeper insights into real-world applications (e.g. highly compressed data).
{credits}
7.0.1 Acknowledgements
This study has been partially supported by SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU.
[…]
The authors have no competing interests.